OpenStudy (anonymous):

Andrew wants to place grouping symbols in the expression 15 + 25 – 18 ÷ 2 so that it has a value of 18.5. Which shows the correct placing of the grouping symbols? A.(15 + 25) – 18 ÷ 2 B.15 + (25 – 18 ÷ 2) C.15 + (25 – 18) ÷ 2 D.15 + 25 – (18 ÷ 2)

OpenStudy (anonymous):

@Willie579

OpenStudy (willie579):

Well this seems pretty easy. I'll try to figure it out. c:

OpenStudy (anonymous):

ok

OpenStudy (willie579):

What the question is saying is find out how to group it where it 's value is 18.5. So let's go through the possible answers and answer them to see if we get 18.5.

OpenStudy (willie579):

For A. (15 + 25) – 18 ÷ 2 (40) - 18 ÷ 2 40 - 18 ÷ 2 40 - 9 =31

OpenStudy (willie579):

So A is incorrect.

OpenStudy (willie579):

On to B.

OpenStudy (anonymous):

ok hehe

OpenStudy (willie579):

15 + (25 – 18 ÷ 2) 15 + (25 - 9) 15 + (16) 15 + 16 =31

OpenStudy (willie579):

NOTE: To do this you must use the PEMDAS rule.

OpenStudy (anonymous):

ok got you

OpenStudy (willie579):

Parenthesis, Exponents, Multiplication*, Division*, Addition**, Subtraction** (Please excuse my dear aunt Sallie!) *Multiplication or Division which ever comes first in the problem ** Addition or Subtraction which ever comes first in the problem

OpenStudy (willie579):

So anyways, on to C.

OpenStudy (willie579):

15 + (25 – 18) ÷ 2 15 + (7) ÷ 2 15 + 7 ÷ 2 15 + 3.5 =18.5

OpenStudy (willie579):

So C is the answer.

OpenStudy (anonymous):

yayyyyy

OpenStudy (willie579):

No problem! :)

OpenStudy (anonymous):

i got another a question

OpenStudy (willie579):

Okay, no problem! :)

OpenStudy (willie579):

What's your question?

OpenStudy (willie579):

@WildQueen